if numbers p,q,r are in A.P. then m7p,m7q  and m7r  are in?
 

Grade:

1 Answers

Tulika Sureka
8 years ago
Answer is they are in G.P
p+r=2q (As they are in AP)
Consider the given three term to be in GP
 
m7pm7r=m2*(7q)
m7*(p+r)=m14q
m7*2q=m14q